To test the flow of water through a 0.25 inch circular nozzle, a flow of 60 gallons was recorded in 10 minutes. Find the velocity in feet per second of the flow of water if the stream is as large as the nozzle.
Answer by Alan3354(69443)

To test the flow of water through a 0.25 inch circular nozzle, a flow of 60 gallons was recorded in 10 minutes. Find the velocity in feet per second of the flow of water if the stream is as large as the nozzle.
====================
60 g/10 mine = 6 gal/minute = 6*232 cubic inches/minute = 23.2 cubic in/sec
---
Assuming 0.25 is the diameter:
Area = pi*r^2 =~ 0.04909 sq inches
Flow rate = 23.2 in^3/sec/(0.04909 in^2) =~474.663 in/sec
=~ 39.555 ft/sec
================================
Notice that including the units decreases the chance of making a mistake.